# Quantum Computing for Query Containmentof Conjunctive Queries

### v1

@@ 1. Introduction
The question of whether one query is contained in another is fundamental: it asks whether, for any possible input, the result of one query will always be a subset of the other. Being able to decide query containment opens numerous opportunities in data management, in particular query optimization, but also data integration, semantic caching, or access control. Variations of the problem have been studied in database theory for decades (cf. e.g. (Abiteboul et al., 1995; Maier, 1983; Marcinkowski and Ostropolski-Nalewaja, 2025)), and for various query languages. Yet, due to the high computational cost of deciding query containment (NP-complete for conjunctive queries with set semantics) or outright infeasibility (undecidable for relational algebra), commercial database systems are restricted to using heuristics or rule-based rewrites. In this work, we focus on the seminal problem of query containment for conjunctive queries under set semantics (QC-CQ), which is known to be NP-complete (Chandra and Merlin, 1977). In particular, we investigate the question of whether this problem can be formulated for solving on quantum computers.
@@
Much of recent research in this field focuses on formulating instances of QUBO optimization problems (Hai et al., 2024), where the objective is to minimize polynomials of degree at most two. This enables execution on quantum annealers, which currently provide access to thousands of qubits. Optimization problems involving higher-degree polynomials are typically addressed using the QAOA algorithm (Farhi et al., 2014). QAOA follows an annealing-inspired strategy but targets another hardware architecture: Gate-based quantum processing units (QPUs) can perform universal computations, yet are currently more restricted in the number of available qubits.
@@
    Our problem formulation can be solved on quantum hardware. Specifically, polynomials with degree of up to 2 can be solved directly by quantum annealers (which currently provide qubits in the range of thousands). For polynomials of arbitrary degree, we use QAOA on circuit-based QPUs (given enough qubits).
@@
    For future integration into a query engine, it is vital that our approach to containment checking only errs on the safe side. Otherwise, a query optimizer risks incorrect results. In fact, our approach is guaranteed to never produce any false positives, as all positives come with a verifiable witness.
@@
    We present a 3-stage workflow to decide QC-CQ. During preprocessing, we perform a simplification step, which we prove to be correct. Our experiments show that simplification is highly effective, either by directly answering QC-CQ, or by reducing the degree of the polynomial so that the problem can be answered using annealing, rather than QAOA.
@@ 2.1. Conjunctive Queries & Query Containment
The same query can be expressed in the syntax of tableau queries (a language similar to Query By Example (QBE) (Abiteboul et al., 1995)) via the answer tuple $(Y_{1})$ and the three relations shown to the right in Figure 1.
@@ 2.2. Quantum Computing & Simulated Annealing
This section is written from a non-physicist perspective and aims to provide an overview, giving sufficient information to database researchers for informed use of the concepts from quantum computing considered in this paper. Particular focus is put on quantum computing methods for solving optimization problems of the form
@@
Here, $p\colon\{0,1\}^{n}\rightarrow\mathbb{R}$ is the objective function; it is a polynomial on $n$ binary variables $\overline{x}=(x_{1},\dotsc,x_{n})$ with real-valued coefficients. The search space is the set $C\subseteq\{0,1\}^{n}$, and the goal is to find a tuple $\overline{x}\in C$ such that the value $p(\overline{x})$ is as small as possible. This optimization problem is called unconstrained if $C=\{0,1\}^{n}$, and it is called constrained if $C\subsetneq\{0,1\}^{n}$. We write $p$ as a sum of monomials, with each monomial being of the form $c\cdot\prod_{i=1}^{n}(x_{i})^{j_{i}}$ where $c$ is a real-valued coefficient, and $j_{i}$ is a non-negative integer exponent associated with variable $x_{i}$ (and, by convention, $(x_{i})^{0}=1$). The degree of such a monomial is defined as the sum $\sum_{i=1}^{n}{j_{i}}$, and the degree of $p$ is the maximum degree of its monomials. Since each variable $x_{i}$ can be assigned only the values 0 and 1, and $0\cdot 0=0$ and $1\cdot 1=1$, any exponent $j_{i}\geqslant 1$ can be replaced with any positive integer $j^{\prime}_{i}$ without changing the function values of the polynomial. If the degree of $p$ is $\leqslant 2$, this implies that we can assume without loss of generality that every monomial of $p$ has degree exactly 2; this setting is known as Quadratic Unconstrained Binary Optimization problems (for short: QUBOs), cf. e.g. (Hai et al., 2024).
Quantum Approximate Optimization Algorithm for gate-based quantum computers. QAOA (Farhi et al., 2014) is a hybrid quantum-classical algorithm designed for combinatorial optimization. It iteratively evaluates the objective function’s expected value for a parameter-dependent quantum state on a quantum processor, using classical feedback to update parameters and converge towards an optimal state. The quantum circuit consists of several layers that transition the system from the ground state of a so-called mixing Hamiltonian (that describes the initial state) to that of the so-called problem Hamiltonian (that describes the objective function). The problem solution is found by sampling the final state. Effectiveness is enhanced by strategies for incorporating problem constraints into the QAOA quantum circuit design (Bärtschi and Eidenbenz, 2020; Zander et al., 2024). The idea is that only the subset of feasible solutions is explored, and the reduced size of the search space improves results.
@@
The complexity of the quantum circuit depends on the specific problem constraints and the objective function $p$. For an arbitrary quadratic polynomial $p$ of $n$ binary variables $\overline{x}=(x_{1},\ldots,x_{n})$ and the constraints considered in this work, the circuits for QAOA with $\ell$ layers have a favorable scaling of $\mathcal{O}(\ell\cdot n^{2})$ gates, $\mathcal{O}(\ell\cdot n)$ depth (runtime), and $\mathcal{O}(n)$ qubits for both the constrained and unconstrained version (Bärtschi and Eidenbenz, 2020; Zander et al., 2024). In general, for a fixed degree $d\leqslant n/2$ square-free polynomial $p$ of $n$ binary variables, the maximum number of monomials is $\mathcal{O}(n^{d})$ (practical instances are typically much sparser.). A degree $d$ monomial requires $\mathcal{O}(d)$ gates and circuit depth, and only a single ancilla qubit (Khattar and Gidney, 2025). It follows that the circuits for QAOA with $\ell$ layers require $\mathcal{O}(\ell\cdot d\cdot n^{d})$ gates and depth, and $\mathcal{O}(n)$ qubits, which is a polynomial scaling in the number of variables $n$. Therefore, they can be efficiently executed on a quantum computer.
@@
Simulated annealing and quantum annealing. After having described a gate-based solver for the discrete optimization problem (1) in the previous paragraphs, we now present the foundations of two further optimization methods that can be used to solve a QUBO, i.e., problem (1) in the unconstrained case where all monomials of $p$ have degree 2. These methods are referred to as simulated annealing and quantum annealing. The term “annealing” is inspired by a cooling technique from metallurgy. Thereby heated metal is cooled in a controlled way such that its atoms have sufficient time to arrange themselves and form stable crystals. This results in a low-energy state of the metal that is considered to be stable. Both simulated annealing and quantum annealing are heuristic search methods moving in a controlled way over a solution landscape for a sufficiently long time until they reach a stable and minimizing state. Due to this feature, both methods contain the term “annealing” in their names. In this work, we use three different solvers provided by the D-Wave library (DWa, 2026), which are briefly described in the following.
@@
Simulated annealing. This solver emulates a controlled cooling procedure by considering a parameter $\beta$, which is related to a possible temperature $T$ of a cooling process as follows: $\beta=1/\left(k_{b}\cdot T\right)$, where $k_{b}$ denotes the Boltzmann constant. The cooling process is defined by a certain range for $\beta$ limited by some $\beta_{\min}$ and $\beta_{\max}$ i.e., $\beta_{\min}\leqslant\beta\leqslant\beta_{\max}$. Typically, one chooses for $\beta_{\min}$ a small value to start at a relatively high temperature, while $\beta_{\max}$ is large corresponding to a low temperature. The parameter $\beta$ influences the probability of circumventing a barrier in the solution landscape between two local minima. As $\beta$ is increasing or $T$ is decreasing, the probability of leaving the current position is decreasing, such that at the end of the process the solver terminates at a possible global minimum. Besides $\beta$, the difference between the values of two neighboring minima influences the probability of moving from the current minimum to the other minimum. In case of large positive differences with respect to $k_{b}{\cdot}T$ or $1/\beta$, it is less likely that the solver leaves the current minimum. Otherwise, if there are negative differences, i.e. if the value of the neighboring minmum is lower, the current minimum will be left and the neighboring local minimum will be attained (cf., (McGeoch, 2022, Sec. 3.1)).
@@
Quantum annealing. Contrary to simulated annealing, quantum annealing does not emulate a cooling process, but a controlled transfer from a quantum state that can be easily prepared to a quantum state yielding with a high probability a global minimizer within a given solution landscape. This is done by introducing an annealing time $t$ and a range for $t$ starting from 0 and ending at some final time $t_{f}$. For a certain annealing time $t$, the quantum state is a weighted sum of the initial state and the final state. The weights in this sum are adapted such that at the end of the annealing process the final state is overweighted compared to the initial state (cf., (McGeoch, 2022, Sec. 3.2 & Chap. 4) and (Rajak et al., 2023)).
@@
Simulated quantum annealing. This solver is a classical version of quantum annealing, i.e. it emulates the behavior of a quantum annealer on a classical computer. The weights for the initial state and the final state are again determined by a parameter $\beta$, which is chosen in a similar way as in case of simulated annealing (Crosson and Harrow, 2016).
@@
Since only a finite number of betas and annealing times can be considered, the user has to provide a parameter denoted by num_sweeps (number of sweeps). As a result, the range for $\beta$ and the annealing time are discretized by num_sweeps different values that are used within the annealing process. In addition to that, a schedule can be chosen to determine how $\beta$ or the annealing time can be changed during an annealing process — either in a linear, or a geometric or a user-defined way. Since heuristic solution methods may not find a global minimizer after a single run, it might be required to run the solution method for multiple times. In D-Wave, one can fix the number of runs by providing an appropriate value for num_reads (number of reads). In terms of the solvers discussed above, a single run is equivalent to a single annealing process.
More background on quantum computing, QAOA, simulated annealing and quantum annealing can be found in Appendix A.
@@ 3.2. Generic choice of $(p,d,C)$
In our generic choice of $(p,d,C)$ we let $C=\{0,1\}^{B_{2}\times A_{1}}$. Our goal is to also define $d$ and $p$ in such a way that Equivalence (2) holds. Here, $p$ will be an integer-valued polynomial on the binary variables $\overline{x}=(x_{i,j})_{(i,j)\in B_{2}\times A_{1}}$; we will define $p$ in such a way, that for an assignment $\overline{x}\in\{0,1\}^{B_{2}\times A_{1}}$ of the binary variables with actual bits, the following is true: the value $p(\overline{x})$ is minimal and equal to $d$ $\iff$ the extended bit-matrix $\overline{x}^{\textup{(e)}}$ represents a function $h\colon A_{2}\to A_{1}$ that, in fact, is a homomorphism from $q_{2}$ to $q_{1}$.
@@
In Sections 4 & 5 we present a quantum computing solution for finding $\overline{x}^{*}$. The size of the search space is $|C_{\textit{gen}}|=2^{n_{2}\cdot n_{1}}$ for $n_{2}:=|B_{2}|$ and $n_{1}:=|A_{1}|$. The number of qubits (and thus, the computational resources) required for our solution is closely related to the number of binary variables (i.e., variables which can be assigned 0-1-values) present in the polynomial $p_{\textit{gen}}(\overline{x})$. The number of binary variables occurring in $p_{\textit{gen}}(\overline{x})$ is $n_{2}{\cdot}n_{1}=|B_{2}|{\cdot}|A_{1}|$. The next two subsections present optional modifications that may decrease the number of binary variables and the size of the search space.
@@ 3.3. Simplification Step
where $(u_{1},\ldots,u_{r})=u$. Here, $x^{\textup{(e)}}_{i,j}$ is the binary variable $x_{i,j}$ if $i\in B_{2}$, and it is a fixed 0-1-value if $i\in A_{2}\setminus B_{2}$. In particular, if there is a $k$ with $u_{k}\in A_{2}\setminus B_{2}$ and $x^{\textup{(e)}}_{u_{k},w_{k}}=0$, then the entire product $\prod_{k=1}^{r}x^{\textup{(e)}}_{u_{k},w_{k}}$ evaluates to 0, and hence the tuple $(w_{1},\ldots,w_{r})\in\textbf{T}_{1}(R)$ has zero contribution in the polynomial $p_{R,u}(\overline{x})$.
@@
$p^{\text{(s)}}_{\textup{(3)}}(\overline{x})\ \ :=\ \ \sum_{R\in\textbf{S}}\ \sum_{u\in\textbf{T}_{2}(R)}\sum_{(w_{1},\ldots,w_{r})\in\textbf{T}_{1}(R)}-\prod_{k=1}^{r}x^{\textup{(s)}}_{u_{k},w_{k}}$
and $p^{\text{(s)}}_{\textup{fct}}(\overline{x}):=\sum_{i\in B^{\text{(s)}}_{2}}\sum_{j,j^{\prime}\in A_{1}\atop j<j^{\prime}}\big(\,x_{i,j}\cdot x_{i,j^{\prime}}\,\big)$. Using Theorem 3.9, the construction of $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$ immediately yields:
@@
During the simplification step, we initialize $B^{\text{(s)}}_{2}$ to $B_{2}$ and then remove from $B^{\text{(s)}}_{2}$ the variables $X_{2},Z_{2}$ and let $x^{\textup{(s)}}_{X_{2},X_{1}}=1$, $x^{\textup{(s)}}_{Z_{2},Z_{1}}=1$, $x^{\textup{(s)}}_{X_{2},j}=0$ for all $j\in A_{1}\setminus\{X_{1}\}$ and $x^{\textup{(s)}}_{Z_{2},j}=0$ for all $j\in A_{1}\setminus\{Z_{1}\}$. Afterwards, we remove $W_{2}$ from $B^{\text{(s)}}_{2}$ and let $x^{\textup{(s)}}_{W_{2},\textsf{\textquotesingle actor\textquotesingle}}=1$ and $x^{\textup{(s)}}_{W_{2},j}=0$ for all $j\in A_{1}\setminus\{\textsf{\textquotesingle actor\textquotesingle}\}$. Then, $B^{\text{(s)}}_{2}$$\emptyset$ and
$p^{\text{(s)}}_{\textup{(3)}}(\overline{x})\ \ =\ \ -\ x^{\textup{(s)}}_{X_{2},X_{1}}\cdot x^{\textup{(s)}}_{Z_{2},Z_{1}}\ -\ x^{\textup{(s)}}_{X_{2},X_{1}}\cdot x^{\textup{(s)}}_{W_{2},\textsf{\textquotesingle actor\textquotesingle}}\ \ =\ \ -2$
and $p^{\text{(s)}}_{\textup{fct}}(\overline{x})=0$. In summary, we obtain: $p_{\textit{sim}}(\overline{x})=-2$, ${d}_{\textit{sim}}=-|\textbf{T}_{2}|=-2$, and $C_{\textit{sim}}=\{0,1\}^{\emptyset\times A_{1}}$ which, by convention, is the set $\{()\}$ consisting only of the empty tuple $()$. We will use this in Example 3.15.

### v2

@@ Abstract.
The source code, data, and/or other artifacts have been made available at https://github.com/miniHive/QC4QCoCQ.
@@ 1. Introduction
The question of whether one query is contained in another is fundamental: it asks whether, for any possible input, the result of one query will always be a subset of the other. Being able to decide query containment opens numerous opportunities in data management, in particular query optimization, but also data integration, semantic caching, or access control. Variations of the problem have been studied in database theory for decades (cf. e.g. (Abiteboul et al., 1995; Maier, 1983; Marcinkowski and Ostropolski-Nalewaja, 2025)), and for various query languages. Yet, due to the high computational cost of deciding query containment (NP-complete for conjunctive queries with set semantics) or outright infeasibility (undecidable for relational algebra), commercial database systems are restricted to using heuristics or rule-based rewrites. In this work, we focus on the query containment problem for conjunctive queries under set semantics (QC-CQ(Chandra and Merlin, 1977). In particular, we investigate whether this problem can be formulated for solving on quantum computers.
@@
Much of recent research in this field focuses on formulating instances of optimization problems in QUBO form (Hai et al., 2024), where the objective is to minimize polynomials of degree at most two. This enables execution on quantum annealers, which currently provide access to thousands of qubits. Optimization problems involving higher-degree polynomials are typically addressed using the QAOA algorithm (Farhi et al., 2014). QAOA follows an annealing-inspired strategy but targets another hardware architecture: gate-based quantum processing units (QPUs), which can perform universal computations, yet are currently more restricted in the number of available qubits.
@@
    Our problem formulation can be solved on quantum hardware. Specifically, polynomials with degree of up to 2 can be directly mapped to quantum annealers (which currently provide qubits in the range of thousands). For polynomials of arbitrary degree, we use QAOA on circuit-based QPUs (given enough qubits).
@@
    For future integration into a query engine, it is vital that our approach to containment checking only errs on the safe side. Otherwise, a query optimizer risks producing incorrect results. In fact, our approach is guaranteed to never produce any false positives, as all positives come with a verifiable witness.
@@
    We present a 3-stage workflow to decide QC-CQ. During preprocessing, we perform a simplification step, which we prove to be correct. Our experiments show that simplification is highly effective: by reducing the number of variables and the degree of the polynomial, the problem can be solved more efficiently.
@@ 2.1. Conjunctive Queries & Query Containment
The same query can be expressed in the syntax of tableau queries (a language similar to Query By Example (QBE) (Abiteboul et al., 1995)) via the answer tuple $(Y_{1})$ and the three relations shown on the right in Figure 1.
@@ 2.2. Quantum Computing & Simulated Annealing
This section is written from a non-physicist perspective and aims to provide an overview, giving sufficient information to database researchers for informed use of the concepts from quantum computing considered in this paper; more background can be found in Appendix A. Particular focus is put on quantum computing methods for solving optimization problems of the form
@@
Here, $p\colon\{0,1\}^{n}\rightarrow\mathbb{R}$ is the objective function; it is a polynomial on $n$ binary variables $\overline{x}=(x_{1},\dotsc,x_{n})$ with real-valued coefficients. The search space is the set $C\subseteq\{0,1\}^{n}$. The goal is to find a tuple $\overline{x}\in C$ such that the value $p(\overline{x})$ is as small as possible. This optimization problem is called unconstrained if $C=\{0,1\}^{n}$ and constrained if $C\subsetneq\{0,1\}^{n}$. We write $p$ as a sum of monomials, with each monomial being of the form $c\cdot\prod_{i=1}^{n}(x_{i})^{j_{i}}$ where $c$ is a real-valued coefficient, and $j_{i}$ is a non-negative integer exponent associated with variable $x_{i}$ (and, by convention, $(x_{i})^{0}=1$). The degree of such a monomial is defined as the sum $\sum_{i=1}^{n}{j_{i}}$, and the degree of $p$ is the maximum degree of its monomials. Since each variable $x_{i}$ can only be assigned the values 0 and 1, and $0\cdot 0=0$ and $1\cdot 1=1$, any exponent $j_{i}\geqslant 1$ can be replaced with any positive integer $j^{\prime}_{i}$ without changing the function values of the polynomial. If the degree of $p$ is $\leqslant 2$, we can assume without loss of generality that every monomial of $p$ has degree exactly 2; this setting is known as Quadratic Unconstrained Binary Optimization problems (for short: QUBOs), cf. e.g. (Hai et al., 2024).
Quantum Approximate Optimization Algorithm for gate-based quantum computers. QAOA (Farhi et al., 2014) is a hybrid quantum-classical algorithm designed for combinatorial optimization. It iteratively evaluates the objective function’s expected value for a parameter-dependent quantum state on a quantum processor, using classical feedback to update parameters and converge towards an optimal state. The quantum circuit consists of several layers that transition the system from the ground state of a so-called mixing Hamiltonian (describes the initial state) to that of the so-called problem Hamiltonian (describes the objective function). The problem solution is found by sampling the final state. Effectiveness is enhanced by incorporating problem constraints into the QAOA quantum circuit design (Bärtschi and Eidenbenz, 2020; Zander et al., 2024). Then, only a subset of feasible solutions is explored, and the reduced size of the search space improves results.
@@
The complexity of the quantum circuit depends on the specific problem constraints and the objective function $p$. In the following, we argue that the complexity scales polynomially in the number $n$ of variables. This indicates that the problem can be efficiently executed on a quantum computer: For an arbitrary quadratic polynomial $p$ of $n$ binary variables $\overline{x}=(x_{1},\ldots,x_{n})$ and the constraints considered in this work, the circuits for QAOA with $\ell$ layers have a favorable scaling of $\mathcal{O}(\ell\cdot n^{2})$ gates, $\mathcal{O}(\ell\cdot n)$ depth (runtime), and $\mathcal{O}(n)$ qubits for both the constrained and unconstrained version (Bärtschi and Eidenbenz, 2020; Zander et al., 2024). In general, for a fixed degree $d\leqslant n/2$ square-free polynomial $p$ of $n$ binary variables, the maximum number of monomials is $\mathcal{O}(n^{d})$ (practical instances are typically much sparser.). A degree $d$ monomial requires $\mathcal{O}(d)$ gates and circuit depth, and only a single ancilla qubit (Khattar and Gidney, 2025). It follows that the circuits for QAOA with $\ell$ layers require $\mathcal{O}(\ell\cdot d\cdot n^{d})$ gates and depth, and $\mathcal{O}(n)$ qubits, which is a polynomial scaling in the number of variables $n$. Hence, they can be efficiently executed on a quantum computer.
@@
Simulated annealing and quantum annealing. Next, we present the foundations of two further methods that can be used to solve a QUBO, i.e., problem (1) (see above) in the unconstrained case where all monomials of $p$ have degree 2. These methods are referred to as simulated annealing and quantum annealing. The term “annealing” is inspired by a cooling technique from metallurgy: heated metal is cooled in a controlled way such that its atoms have sufficient time to arrange themselves and form stable crystals. This results in a low-energy state of the metal that is considered to be stable. Both simulated annealing and quantum annealing are heuristic search methods moving in a controlled way over a so-called solution landscape or energy landscape (i.e., possible assignments of the binary variables; henceforth these are called states) for a sufficiently long time until they reach a stable and minimizing (w.r.t. the polynomial value) state. Thus, both methods contain the term “annealing” in their names. A minimum state corresponds to an optimal solution. In this work, we use three different solvers provided by the D-Wave library (DWa, 2026), which are briefly described in the following.
@@
Simulated annealing. This solver emulates a controlled cooling procedure by considering a parameter $\beta$, which is related to a possible temperature $T$ of a cooling process via $\beta=1/\left(k_{b}\cdot T\right)$, where $k_{b}$ denotes the Boltzmann constant. The cooling process is defined by a certain range for $\beta$ limited by some $\beta_{\text{start}}$ and $\beta_{\text{end}}$ with $\beta_{\text{start}}\leqslant\beta\leqslant\beta_{\text{end}}$. Typically, one chooses for $\beta_{\text{start}}$ a small value to start at a relatively high temperature, while $\beta_{\text{end}}$ is large corresponding to a low temperature; $\beta$ influences the probability of circumventing a barrier (value difference) in the energy landscape between neighboring states (i.e., states that differ in one bit). As $\beta$ increases (i.e., $T$ decreases), the probability of leaving a lower position decreases, such that the solver becomes less exploratory with time and finally terminates at a possible global minimum. Besides $\beta$, the difference between the values of neighboring states influences the probability of moving between them. In case of large positive differences w.r.t. $1/\beta$, it is less likely that the solver leaves a current state. Otherwise, if there are negative differences, i.e. if the value of the neighboring state is lower, the current state will be left and the lower-valued neighbor will always be attained (cf., (McGeoch, 2022, Sec. 3.1)).
@@
Quantum annealing. Contrary to simulated annealing, quantum annealing does not emulate a cooling process, but rather a controlled transfer from a quantum state that can be easily prepared to a quantum state that yields with high probability a global minimizer within a given energy landscape. This is done by introducing an annealing time $t$ that ranges from 0 to some final time $t_{f}$. For a certain annealing time $t$, the quantum state is a weighted sum of the initial state and the final state. The weights in this sum are slowly adapted such that at the end of the annealing process, the final state is overweighted compared to the initial state (cf., (McGeoch, 2022, Sec. 3.2 & Chap. 4) and (Rajak et al., 2023)). In contrast to simulated annealing, the probability of circumventing energy barriers for quantum annealing only depends on the width of the barrier (distance between states) rather than the height (difference between values).
@@
Simulated quantum annealing. This solver is a classical version of quantum annealing, i.e. it emulates the behavior of a quantum annealer on a classical computer. The weights for the initial state and the final state are again determined by a parameter $\beta$, which is chosen in a similar way as in the case of simulated annealing (Crosson and Harrow, 2016).
@@
Since only a finite number of betas and annealing times can be considered, the user has to provide a parameter denoted by num_sweeps (number of sweeps). As a result, the range for $\beta$ and the annealing time are discretized by num_sweeps different values that are used within the annealing process. In addition to that, a schedule can be chosen to determine how $\beta$ or the annealing time can be changed during an annealing process — either in a linear, or a geometric or a user-defined way. Since heuristic solution methods may not find a global minimizer after a single run, it might be required to run the solution method for multiple times. In D-Wave, one can fix the number of runs by providing an appropriate value for num_reads (number of reads). In terms of the solvers discussed above, a single run/read is equivalent to a single annealing process.
@@ 3.2. Generic choice of $(p,d,C)$
In our generic choice of $(p,d,C)$ we let $C=\{0,1\}^{B_{2}\times A_{1}}$. Our goal is to define $d$ and $p$ in such a way that Equivalence (2) holds. Here, $p$ will be an integer-valued polynomial on the binary variables $\overline{x}=(x_{i,j})_{(i,j)\in B_{2}\times A_{1}}$; we will define $p$ in such a way, that for an assignment $\overline{x}\in\{0,1\}^{B_{2}\times A_{1}}$ of the binary variables with actual bits, the following is true: the value $p(\overline{x})$ is minimal and equal to $d$ $\iff$ the extended bit-matrix $\overline{x}^{\textup{(e)}}$ represents a function $h\colon A_{2}\to A_{1}$ that, in fact, is a homomorphism from $q_{2}$ to $q_{1}$.
@@
In Sections 4 & 5 we present a quantum computing solution for finding $\overline{x}^{*}$. The size of the search space is $|C_{\textit{gen}}|=2^{n_{2}\cdot n_{1}}$ for $n_{2}:=|B_{2}|$ and $n_{1}:=|A_{1}|$. The number of qubits (and thus, the computational resources) required for our solution is closely related to the number of binary variables present in the polynomial $p_{\textit{gen}}(\overline{x})$. The number of binary variables occurring in $p_{\textit{gen}}(\overline{x})$ is $n_{2}{\cdot}n_{1}=|B_{2}|{\cdot}|A_{1}|$. The next two subsections present optional modifications that may decrease the number of binary variables and the size of the search space.
@@ 3.3. Simplification Step
where $(u_{1},\ldots,u_{r})=u$. Here, $x^{\textup{(e)}}_{u_{k},w_{k}}$ is the binary variable $x_{u_{k},w_{k}}$ if $u_{k}\in B_{2}$, and it is a fixed 0-1-value if $u_{k}\in A_{2}\setminus B_{2}$. In particular, if there is a $k$ with $u_{k}\in A_{2}\setminus B_{2}$ and $x^{\textup{(e)}}_{u_{k},w_{k}}=0$, then the entire product $\prod_{k=1}^{r}x^{\textup{(e)}}_{u_{k},w_{k}}$ evaluates to 0, and hence the tuple $(w_{1},\ldots,w_{r})\in\textbf{T}_{1}(R)$ has zero contribution in the polynomial $p_{R,u}(\overline{x})$.
@@
During simplification, we initialize $B^{\text{(s)}}_{2}$ to $B_{2}$ and then remove from $B^{\text{(s)}}_{2}$ the variables $X_{2},Z_{2}$, and let $x^{\textup{(s)}}_{X_{2},X_{1}}=1$, $x^{\textup{(s)}}_{Z_{2},Z_{1}}=1$, $x^{\textup{(s)}}_{X_{2},j}=0$ for all $j\in A_{1}\setminus\{X_{1}\}$ and $x^{\textup{(s)}}_{Z_{2},j}=0$ for all $j\in A_{1}\setminus\{Z_{1}\}$. Afterwards, we remove $W_{2}$ from $B^{\text{(s)}}_{2}$ and let $x^{\textup{(s)}}_{W_{2},\textsf{\textquotesingle actor\textquotesingle}}=1$ and $x^{\textup{(s)}}_{W_{2},j}=0$ for all $j\in A_{1}\setminus\{\textsf{\textquotesingle actor\textquotesingle}\}$. Then, $B^{\text{(s)}}_{2}$$\emptyset$ and $p^{\text{(s)}}_{\textup{(3)}}(\overline{x})=-x^{\textup{(s)}}_{X_{2},X_{1}}{\cdot}x^{\textup{(s)}}_{Z_{2},Z_{1}}-x^{\textup{(s)}}_{X_{2},X_{1}}{\cdot}x^{\textup{(s)}}_{W_{2},\textsf{\textquotesingle actor\textquotesingle}}=-2$ and $p^{\text{(s)}}_{\textup{fct}}(\overline{x})=0$. Thus, we obtain: $p_{\textit{sim}}(\overline{x})=-2$, ${d}_{\textit{sim}}=-|\textbf{T}_{2}|=-2$, and $C_{\textit{sim}}=\{0,1\}^{\emptyset\times A_{1}}$ which, by convention, is the set $\{()\}$ consisting only of the empty tuple $()$. We leverage this in Example 3.15.